OpenStudy (anonymous):

find the domain of (x+2)(1/x+2)

OpenStudy (campbell_st):

expand 1 + 2x + 2/x + 4 or 5 +2x + 1/x domain is all real x except x cannot equal zero.

OpenStudy (anonymous):

well first, the function is a polynomial. So the initial domain of all polynomials is all real numbers \[(-\infty, \infty) \] Now we have to figure out where the function is undefined. We know that if x=-2, then you get: (0) * (1/0) And 1/0 is undefined. So the function can never take on the value of x=-2 which means your domain is simply: \[(-\infty, -2)\cup(-2, \infty)\]